OpenStudy (anonymous):

Simplify the following expression. Write the result using positive exponents only. 8x^-1

jhonyy9 (jhonyy9):

do you know the property of exponents ? a^-1 = ?

OpenStudy (anonymous):

no 8x^-1 was the only information given.

jhonyy9 (jhonyy9):

so for your help a^-1 = 1/a can us it for your exercise ?

jhonyy9 (jhonyy9):

so this mean that x^-1 = 1/x yes ?

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

so than there is 8*(1/x) = ???

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

do you understand it now ?

OpenStudy (anonymous):

yes thank you

jhonyy9 (jhonyy9):

ok good luck bye

jhonyy9 (jhonyy9):

can you writ here what will be the result ?